OpenStudy (anonymous):

What is the slope of a line with the equation y-2=-3/4(x+5)

OpenStudy (anonymous):

re-write the equation in the form: y = mx + b The slope will be the m. Or you can recognize that the equation is in the point-slope form: y-y1 = m(x-x1) And the slope is once again the m.

OpenStudy (anonymous):

Its -3/4

OpenStudy (anonymous):

y=-3/4x-15/4+8/4 y=-3/4x-7/4 slope, m=?